The Detonation Database Project

Explosion Dynamics Laboratory

Graduate Aeronautical Laboratories California Institute of Technology

Michael Kaneshige · Joseph E. Shepherd · Andrzej Teodorczyk

Introduction

Background Experimental data are particularly important for studies and analyses of gaseous detonations. This is a reflection of the largely empirical nature of our knowledge about detonation behavior, a consequence of the complexity of detonation phenomena, which encompass aspects of compressible fluid dynamics, thermodynamics, and gas phase chemistry. The large number of publications containing detonation data presents a formidable challenge to detonation researchers wishing to utilize previous results.

Purpose This database was initially started for the purposes of supporting our activities at Caltech in cell size measurement and correlation with ZND models based on detailed chemical reaction mechanisms. It quickly became clear that it was also very useful for other experimental or numerical studies and explosion hazard analyses carried out in the Explosion Dynamics Laboratories at the California Institute of Technology.

Accessibility The project has evolved into an electronic publication http://www.galcit.caltech.edu/˜ jeshep/detn db/html/

and a print-on-paper version, available as a PostScript file (1.63 MB) that can be downloaded through the WWW version. Each format contains the same information with somewhat different features. For instance, the printed document is more readable, but the WWW version offers a search engine.

Contents To keep the diversity of data manageable, the scope is limited to several canonical categories, namely cell size, critical tube diameter, critical energy, and minimum tube diameter, but the types of fuels, oxidizers, diluents, and conditions is unlimited.

Statistics The usefulness of the database depends on its convenience and its coverage. The database cites 120 references and is organized into 281 separate data sets. These data sets are combined and summarized in 78 graphs. Altogether, this amounts to 185 pages of printed matter, with explanatory text. The success of the database is best measured by its popularity. During the month of July, 1997, the World Wide Web version received a total of 3490 hits, from 103 separate sources.

Organization

For the sake of presentation, the database is composed of a list of references, a collection of data sets (tables), and a collection of graphs that combine and display the tabulated data. These are cross referenced as shown below, so that data that appear useful from a graph can be accessed directly as tabulated numbers and then traced to the original source. The document exists in printed and electronic (WWW) form. The electronic form incorporates a search engine that interactively organizes links to the references, data sets, and graphs that match the search parameters.

Currently, the data consist of four categories: cell size, critical tube diameter, minimum tube diameter, and critical energy. Within each category, the data are generally grouped by fuel. Naturally, some exceptions are necessary because important data presented in the literature do not always conform to this categorization.

Function and Operation

The database document is never complete, but grows as new data are located and added. With respect to the printed form and the static WWW form, this only means that the various graphs, references, and links must be updated periodically. The search engine is of course dynamic in nature, returning all parts of the latest static document that match the search parameters.

The processes and tools used to update the document in all its forms, including the dynamically generated search reports, are shown in the diagram above. The flowchart is representative of the complexity of the project, although the actual update procedure is heavily automated.

Contributing Data

Our efforts to locate and digitize important data sets and integrate them into the database have been, by no means, exhaustive. If you have a data set or are aware of a reference that has not been included in the database but would contribute to its mission, please bring it to our attention. We are only interested in data that appear in citable documents, and peer reviewed archival literature sources are preferred. If the source is difficult to access, we will need copies of it. Send contributions to either of the following addresses. Email and ASCII files are the preferred modes of communication.

Professor Joseph E. Shepherd

MC 205-45

Graduate Aeronautical Laboratories California Institute of Technology Pasadena, California 91125 jeshep@galcit.caltech.edu

If you are sending your own data, tabulated form is preferred. Please include as much information as possible about the conditions of the experiments. Note that the database may, in the future, include more types of information than it currently does, for instance qualitative observations by the experimenter.

A typical data set should be formatted something like (although delimiters other than comma-space can be handled):

#Initial Pressure (atm), Initial Pressure (kPa), Cell Width
0.0945, 9.57285, 18.6476
0.1313, 13.30069, 11.5441
0.1997, 20.22961, 6.4376
0.2611, 26.44943, 4.1347
0.3269, 33.11497, 3.3348
0.3269, 33.11497, 3.1164
0.3923, 39.73999, 2.6078